OpenStudy (anonymous):

can someone please help me, I have difficulty understanding math. for example (2squareroot(5)+2)^2

OpenStudy (anonymous):

\[(2\sqrt{5}+2)^2=(2\sqrt{5}+2)(2\sqrt{5}+2)=2\sqrt{5}(2\sqrt{5}+2)+2(2\sqrt{5}+2)\]

OpenStudy (anonymous):

Go from there.

OpenStudy (anonymous):

Foil would work here.

OpenStudy (anonymous):

thanks a bunch the thing is that I dont know how to work with the square roots :(

OpenStudy (anonymous):

So...\[\left( 2\sqrt{5} + 2 \right)\left( 2\sqrt{5} + 2 \right) = 2\sqrt{5}(2\sqrt{5}) + 2\sqrt{5}(2) + 2(2\sqrt{5}) + 2(2)\]

OpenStudy (anonymous):

well, if you do \[(2\sqrt{5})(2\sqrt{5})\] it's 4*5 because the square roots cancel out because they're squared

OpenStudy (anonymous):

It's not different from regular numbers, really. Take the first term as an example: \[2\sqrt{5}(2\sqrt{5})=4\sqrt{5}\sqrt{5}=4\cdot 5=20\]

OpenStudy (anonymous):

The rest of your problem reduces to things you already know, I think.

OpenStudy (anonymous):

and \[2(2\sqrt{5}) = 4\sqrt{5}\]

OpenStudy (anonymous):

I know i may be asking for much but It's been a long while i haven't gone to school and ive been having a hard time and i really don' understand it

OpenStudy (anonymous):

it's like the square root 5 is it's own variable, 2x * 2 = 4x but if you square it, then it turns into a normal # - 5

OpenStudy (anonymous):

Excellent analogy. That is usually how it is explained. It also has to do with the fact that you can't "simplify" two numbers which do not have the same exponent. For example, \(2\sqrt{2}\) cannot be simplified further.

OpenStudy (anonymous):

I appreciate you alls help but i dont get it

OpenStudy (anonymous):

Precisely what is confusing you? Do you see how: \[(2\sqrt{5}+2)^2=(2\sqrt{5}+2)(2\sqrt{5}+2)\] ?

OpenStudy (anonymous):

yes that i do see why its twice because of the exponent 2 and how that becomes 4square root 5 because 2*2=4

OpenStudy (anonymous):

what do you not get?

OpenStudy (anonymous):

but then it has a +2 and what do i do with that one?

OpenStudy (anonymous):

i square it as well?

OpenStudy (anonymous):

oh at the end, 2*2 = 4.

OpenStudy (anonymous):

you treat 2sqrt(5) and 2 as different #s

OpenStudy (anonymous):

Okay, let me work you through it.

OpenStudy (anonymous):

the actual problem is (2sqrt(5) +2)^2 so then my final answer would be what? 4sqrt(5)+4?

OpenStudy (anonymous):

ok thanks alot

OpenStudy (anonymous):

\[(2√5+2)(2√5+2)=2√5(2√5)+2√5(2)+2(2√5)+2(2) = 20 + 4√5 + 4√5 + 4 = 24 + 8√5\]

OpenStudy (anonymous):

oh, you can't see all of it. well then you have 20 + 4√5 + 4√5 + 4 you can combine 4√5 and 4√5 to get 8√5 and 20 + 4 = 24 so your final answer should be 24 + 8√5

OpenStudy (anonymous):

do you understand how i got that?

OpenStudy (anonymous):

no the 20 how did u get it ?

OpenStudy (anonymous):

ive got this other one (6-sqrt(3)) (6+sqrtroot(3)

OpenStudy (anonymous):

2√5 * 2√5 = 2*2*√5*√5 = 4*√25 = 4*5 = 20

OpenStudy (anonymous):

for your other problem, it's called difference of two squares. \[\left( 6 - \sqrt{3} \right)\left( 6 + \sqrt{3}\right) = 36 - 3 = \]

OpenStudy (anonymous):

if the two binomials fit this formula: \[\left(a + b \right) \left( a - b \right)\] then the answer is \[a ^{2} - b ^{2}\]

OpenStudy (anonymous):

this is because the in-between terms would be -ab + ab, so they cancel out. so your other problem's answer is 36 - 3, or 33.